as Revew An object has deraity P each ofl the okjects Eree dmensions is nonased by a factor of 2 without changing the objects mass By what facfor will the derailty change?
Answer #1

A) And: 1

The density will not change.

B) And: 1/8 or 0.125

The density will decrease by a factor of 8.

The volume has increased by 2^3=8 times, but Mass remains the same.

So density_new = m/8V =density_original/8